Find the fourier integral for f(x)= C |x|<=1 and f(x)= 0 |x|> 1 we here C is constant


1
Expert's answer
2022-03-14T17:04:36-0400

Calculate the coefficients:

"A(\lambda)=\frac{1}{\pi}\int_{-\infty}^{+\infty}f(x)\cos(\lambda x)dx=\frac{c}{\pi}\int_{-1}^{+1}\cos(\lambda x)dx\\
=\frac{2c}{\lambda\pi}\sin(\lambda)"

Due to the fact that "f" is an even function "B(\lambda)\equiv0" . So, fourier integral:

"f(x)=\int_0^{+\infty}\frac{2c}{\lambda\pi}\sin(\lambda)\cos(\lambda x)d\lambda"